# Cross-lingual Speech Emotion Recognition: Humans vs. Self-Supervised Models

### v1

@@ III-A Datasets and Models
  • TJD: a non-public Tianjin (TJ) Chinese dialect corpus collected in our previous work [20]. It was recorded and annotated at Tianjin University by two native Tianjin dialect speakers who were university students. It includes three functional categories (question, negation, expectation), approximated to emotions due to high acoustic similarity. According to annotators, negation resembles anger, and expectation resembles happiness. Tianjin dialect is known for its complex tone sandhi patterns while featuring a similar but slightly different tone system than Mandarin [21]. Native speakers of the Tianjin dialect convey emotions more directly with noticeable sonorous vowels and faster speech [22].
For the models, we use three W2V2 base models pre-trained on Mandarin CN [^1] https://huggingface.co/TencentGameMate/chinese-wav2vec2-base, DE [^2] https://huggingface.co/facebook/wav2vec2-base-de-voxpopuli-v2, and EN [^3] https://huggingface.co/facebook/wav2vec2-base-960h, along with a WavLM large model trained on EN emotional speech  [^4] https://huggingface.co/speechbrain/emotion-diarization-wavlm-large.
The following tasks are conducted using different models and datasets for specific purposes.
@@ III-C PEFT of SSL Models for Cross-lingual SER
After the layer-wise analysis, the best-performing layers are further fine-tuned using various PEFT strategies to enhance performance. We apply the Low-Rank Adapter (LoRA) [24], Bottleneck Adapter (BA) [25], and Weighted Gating (WG). Additionally, a two-stage fine-tuning [5] is performed: the model is first fine-tuned on the source language, then on the target language once the first fine-tuning converges.
@@ III-D Comparison of SSL Models with Human Evaluation
Six native DE speakers (one male, five female) and six native CN speakers (two male, four female), with no prior knowledge of each other’s language, are recruited for the human evaluation from the Univ. of Edinburgh and Tianjin Univ. All participants have studied English for many years with sufficient skills (e.g., IELTS score $\geq$ 6.5). The webMUSHRA interface [26] is used to create the experimental tests.
@@ IV-A Experimental Settings
For SER, to reduce the effect of varying training data sizes, we use the same amount of data for CN, DE, and EN. To ensure a balanced emotion distribution, we use an equal number of samples for each emotion. Specifically, for ESD, PAVOQUE, and IEMOCAP, we apply 5-fold cross-validation for model training: 400 utterances per emotion category, totaling 1,600 utterances per dataset, are used for training. Similarly, 200 utterances are randomly selected for validation and test sets, respectively. Given the difficulty of performing human evaluation on all the data, for comparison with human evaluation, we select 12 sentences per emotion category, totaling 144 utterances for each language. The model settings are as follows:

### v2

@@ III-A Datasets and Models
  • IEMOCAP: an English (EN) emotion corpus [19], where five male and female speakers were paired to record scripted and improvised emotional utterances, divided into nine emotion categories.
@@
  • TJD: a non-public Tianjin (TJ) Chinese dialect corpus collected in our previous work [21]. It was recorded and annotated at Tianjin University by two native Tianjin dialect speakers. It includes three functional categories (question, negation, expectation), approximated to emotions due to high acoustic similarity. According to annotators, negation resembles anger, and expectation resembles happiness. Tianjin dialect is known for its complex tone sandhi patterns while featuring a similar but slightly different tone system to Mandarin [22]. Native speakers of the Tianjin dialect convey emotions more directly with noticeable sonorous vowels and faster speech [23].
For the models, we use three W2V2 base models pre-trained on Mandarin CN [^1] https://huggingface.co/TencentGameMate/chinese-wav2vec2-base, DE [^2] https://huggingface.co/facebook/wav2vec2-base-de-voxpopuli-v2, and EN [^3] https://huggingface.co/facebook/wav2vec2-base-960h, along with a WavLM large model trained on EN emotional speech  [^4] https://huggingface.co/speechbrain/emotion-diarization-wavlm-large. The following tasks are conducted using different models and datasets for specific purposes.
@@ III-C PEFT of SSL Models for Cross-Lingual SER
After the layer-wise analysis, the best-performing layers are further fine-tuned using various PEFT strategies to enhance performance. We apply the Low-Rank Adapter (LoRA) [25], Bottleneck Adapter (BA) [26], and Weighted Gating (WG) [5]. Additionally, a two-stage fine-tuning [5] is performed: the model is first fine-tuned on the source language, then on the target language once the first fine-tuning converges.
@@ III-D Comparison of SSL Models with Human Evaluation
Six native DE speakers (one male, five female) and six native CN speakers (two male, four female), with no prior knowledge of each other’s language, are recruited for the human evaluation from the University of Edinburgh and Tianjin University. All participants have studied English for many years with sufficient skills (e.g., IELTS score $\geq$ 6.5). The webMUSHRA interface [27] is used to create the experimental tests.
@@ IV-A Experimental Settings
For SER, to reduce the effect of varying training data sizes, we use the same amount of data for CN, DE, and EN. To ensure a balanced emotion distribution, we use an equal number of samples for each emotion. Specifically, for ESD, PAVOQUE, and IEMOCAP, we apply 5-fold cross-validation for model training: 400 utterances per emotion category, totaling 1,600 utterances per dataset, are used for training. Similarly, 200 utterances are randomly selected for validation and test sets, respectively. Given the difficulty of performing human evaluation on all the data, for comparison with human evaluation, we select 12 sentences per emotion category, totaling 144 utterances for all languages (12 sentences $\times$ 4 emotions $\times$ 3 datasets). The model settings are as follows: